# Yogesh Kulkarni, Pooyan Fazli

### v1

@@ Introduction
We use Qwen2.5-Omni-7B (Xu et al. 2025) as the backbone model and implement AVATAR using the MS-Swift (Zhao et al. 2025b) framework for GRPO training. As shown in Figure 3, we begin with a cold-start SFT phase on general image and video QA datasets, including Video-R1-CoT (Feng et al. 2025) and TPO (Li et al. 2025b). This is followed by three RL stages of increasing complexity: Stage 1 builds foundational visual reasoning using general video QA data from Video-R1 (Feng et al. 2025). Stage 2 incorporates audio-visual QA using AVQA (Yang et al. 2022) and AVE (Tian et al. 2018), with audio cues annotated via Kimi-Audio (KimiTeam et al. 2025). Stage 3 focuses on audio-visual object localization, using AVSBench (Zhou et al. 2022) and stepwise rewards from the InternVL-3B (Zhu et al. 2025) judge. To ensure computational efficiency while preserving temporal information, we uniformly sample 32 video frames during both training and evaluation. We also apply TAS throughout all RL stages. All experiments are conducted on 8 NVIDIA L40S GPUs.
State-of-the-art-models.
We compare AVATAR with a broad range of models across audio-visual and video-only modalities. For audio-visual models, we evaluate against Qwen2.5-Omni-7B (Xu et al. 2025), EchoInk (Xing et al. 2025), Omni-R1 (Zhong et al. 2025), HumanOmni (Zhao et al. 2025a), Ola-7B (Liu et al. 2025b), and AV-Reasoner (Lu et al. 2025). To assess the impact of audio integration, we also compare with SOTA video-only models: Qwen2.5VL (Bai et al. 2025), Video-R1 (Feng et al. 2025), Video-RFT (Wang et al. 2025b), and TW-GRPO (Dang et al. 2025).
@@
For audio-visual tasks, we use OmniBench (Li et al. 2025e) (image-audio-text reasoning), DailyOmni (Zhou, Wang, and Wu 2025) (ambient sound scenarios), AV-Counting (Lu et al. 2025) (cross-modal counting), AV-Odyssey (Gong et al. 2024) (audio-visual alignment), WorldSense (Hong et al. 2025) (omnimodal understanding in-the-wild), and IntentBench (Zhao et al. 2025a) (reasoning about human intent and actions). For video understanding and reasoning, we evaluate on MVBench (Li et al. 2024) (temporal perception-to-cognition QA), Video-MME (Fu et al. 2025) (full-spectrum video comprehension), LVBench (Wang et al. 2024) (long-video reasoning), Video-Holmes (Cheng et al. 2025) (causal inference in suspense clips), MMVU (Zhao et al. 2025c) (domain-expert knowledge analysis), and TOMATO (Shangguan et al. 2025) (temporal reasoning).
Ablation Configuration Audio-Visual Video Reasoning
Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
Curriculum SFT Only 45.8+1.6 45.2+1.2 30.1+0.3 41.8+1.2 62.1+1.9 28.9-0.1
SFT + Stage 1 RL 47.2+3.0 45.8+1.8 30.6+0.8 43.5+2.9 64.2+4.0 29.4+0.4
SFT + Stages 1+2 RL 48.6+4.4 46.7+2.7 31.8+2.0 44.2+3.6 65.1+4.9 30.1+1.1
SFT + Stages 1+2+3 RL (Ours) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
TAS Linear Decay Weights 48.3+4.1 46.2+2.2 31.6+1.8 44.3+3.7 64.8+4.6 30.2+1.2
Linear Incline Weights 48.5+4.3 46.4+2.4 31.7+1.9 44.6+4.0 65.0+4.8 30.4+1.4
No TAS (Uniform Weights) 47.8+3.6 45.9+1.9 31.6+1.8 43.7+3.1 64.8+4.6 29.3+0.3
U-Shaped Weights (Ours) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
Table 4: Ablation studies on curriculum design, reward strategy, and TAS.
Results
@@
Table 1 shows that AVATAR significantly outperforms the Qwen2.5-Omni baseline on audio-visual reasoning tasks, with absolute gains of +4.9 on OmniBench, +3.0 on DailyOmni, and +2.3 on AV-Odyssey. Unlike standard GRPO, which often stalls on overly easy or hard samples (due to vanishing advantages), our stratified replay buffer keeps training in a “challenging but solvable” zone. This mitigates overfitting and builds a more robust reasoning foundation, enabling AVATAR to outperform other GRPO-based methods, surpassing AV-Reasoner by 0.8 and Omni-R1 by 2.2 on OmniBench alignment tasks. TAS further enhances performance on benchmarks that require precise temporal alignment, such as AV-Odyssey. By creating larger, more targeted gradients for the initial tokens that identify an audio cue and the final tokens that synthesize the corresponding visual event, TAS directly reinforces the cross-modal causal links required for such tasks.
This targeted credit assignment enhances AVATAR’s performance on audio-visual tasks like AV-Counting, contributing to stronger generalization across diverse, real-world scenarios, as seen in WorldSense.
@@
As shown in Table 2, AVATAR achieves SOTA performance on general video understanding benchmarks, such as MVBench (66.4) and LVBench (38.4), and excels in complex reasoning tasks. In particular, it delivers significant gains on multi-step causal inference benchmarks, such as Video-Holmes (+4.5) and MMVU (+5.4), enabled by two key components: the stratified replay buffer, which ensures exposure to challenging data, and TAS, which guides reasoning by focusing on the planning and synthesis phases of the token sequence. This is especially evident when compared to other credit assignment methods. TW-GRPO’s statistical token weighting results in unstable performance and even degrades on temporal fine-grained tasks, such as TOMATO. In contrast, TAS leverages the reasoning chain’s temporal structure to provide a more stable learning signal, enabling AVATAR to consistently achieve top performance across reasoning benchmarks.
Ablation Studies
How efficient is AVATAR compared to baseline GRPO?
Figure 4 shows AVATAR’s superior sample efficiency and final performance over the GRPO baseline. GRPO plateaus at a 0.4 reward, solving only simpler queries due to uniform credit assignment. In contrast, AVATAR reaches 0.75 in just 2,500 iterations, a 35% efficiency gain. Its initial dip reflects the off-policy design, where the stratified buffer introduces harder samples. AVATAR quickly adapts via TAS-based credit assignment, converging efficiently while maintaining exploration at high reward levels. This shows its ability to learn from challenging samples that the baseline fails to leverage.
(a) Baseline Method
@@
Figure 4: Comparison of accuracy rewards during training.
What is the effect of each component in AVATAR?
Our ablation study (Table 3) highlights the impact of each component. Adding the replay buffer to standard GRPO yields substantial gains (e.g., +3.6 on OmniBench) by improving data efficiency and generalization. TAS alone also enhances performance (e.g., +4.1 on MMVU) by addressing uniform credit assignment and strengthening reasoning. Combining both further improves results, as the replay buffer enhances sample quality while TAS guides learning. Finally, the hint mechanism provides additional gains (e.g., +0.9 on OmniBench) by helping the model escape local optima, leading to the best overall performance of the full AVATAR system.
What is the effect of the curriculum and TAS shape?
Our ablations in Table 4 validate the effectiveness of both the three-stage training curriculum and the U-shaped TAS weighting. The curriculum shows a clear trend of consistently positive returns. Stage 1 RL yields the largest improvements over SFT (e.g., +2.1 on MMVU) by building core visual reasoning skills. Stage 2 adds further improvements (e.g., +1.4 on OmniBench) as the model learns to handle more complex cross-modal alignment, including audio cues. Stage 3 offers additional gains (e.g., +0.9 on Video-Holmes), reflecting the inherent difficulty of fine-grained localization tasks. The TAS ablation supports our core hypothesis: weighting only the planning (linear decay) or synthesis (linear incline) phases leads to weaker performance. In contrast, the U-shaped weighting, which emphasizes both early planning and late synthesis, proves most effective.
Config Halluc TP AC
Baseline GRPO 35.6 53.6 32.5
AVATAR w/o Off-Policy 40.1+4.5 54.8+1.2 34.1+1.6
AVATAR w/o TAS 48.9+13.3 56.1+2.5 35.8+3.3
AVATAR (Full) 51.2+15.6 57.2+3.6 37.3+4.8
How do AVATAR’s components impact audio-visual reasoning?
Conclusion

### v2

@@ 5 Experiments
We evaluate AVATAR on Qwen2.5-Omni-7B [qwen25omni] and Ola-7B [ola], two open-source MLLMs that provide native audio and video support and are the current SOTA in this space. We implement AVATAR using the MS-Swift [swift] framework for GRPO training. Our training follows the curriculum detailed in Table 1 and Figure 3, beginning with a cold-start SFT phase (S0) followed by three RL stages (S1-S3) of increasing complexity, each targeting different reasoning skills using specific datasets and reward configurations. Stage 2 uses audio cues annotated via Kimi-Audio [kimi], and Stage 3 employs stepwise rewards from an InternVL3 [internvl3] judge. We uniformly sample 32 video frames during both training and evaluation. TAS is applied throughout all RL stages. We attach LoRA [hu2021lora] adapters to the Vision Encoder, Audio Encoder, MLP, and the LLM with parameters $r=16$ and $\alpha_{\text{LoRA}}=32$. The key hyperparameters for AVATAR are $\lambda_{TAS}=0.3,\alpha=0.6,\beta=0.1$.
State-of-the-art Models.
We compare AVATAR with a broad range of models across audio-visual and video-only modalities. For audio-visual models, we evaluate against Qwen2.5-Omni-7B [qwen25omni], EchoInk [echoink], Omni-R1 [omnir1], HumanOmni [humanomni], Ola-7B [ola], and AV-Reasoner [avreasoner]. To assess the impact of audio integration, we also compare with SOTA video-only models: Qwen2.5VL [qwen25vl], Video-R1 [videor1], Video-RFT [videorft], and TW-GRPO [twgrpo].
@@
For audio-visual tasks, we use OmniBench [omnibench] (image-audio-text reasoning), DailyOmni [dailyomni] (ambient sound scenarios), AV-Counting [avreasoner] (cross-modal counting), AV-Odyssey [avodyssey] (audio-visual alignment), WorldSense [worlsense] (omnimodal understanding in-the-wild), and IntentBench [humanomni] (reasoning about human intent and actions). For video understanding and reasoning, we evaluate on MVBench [mvbench] (temporal perception-to-cognition QA), Video-MME [videomme] (full-spectrum video comprehension), LVBench [lvbench] (long-video reasoning), Video-Holmes [videoholmes] (causal inference in suspense clips), MMVU [mmvu] (domain-expert knowledge analysis), and TOMATO [tomato] (temporal reasoning).
Table 5: Ablation studies on training curriculum and advantage shaping strategies. TAS significantly outperforms uniform weights and inverse parabolic weights, validating our hypothesis that early and late reasoning phases are most critical.
Ablation Configuration Audio-Visual Video Reasoning Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0 Training Curriculum SFT Only 45.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.6})}$ 45.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 30.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ 41.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 62.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 28.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{-0.1})}$ SFT + Stage 1 RL 47.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 45.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 30.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.8})}$ 43.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.9})}$ 64.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.0})}$ 29.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.4})}$ SFT + Stages 1+2 RL 48.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.4})}$ 46.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.7})}$ 31.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.0})}$ 44.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.6})}$ 65.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 30.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.1})}$ SFT + Stages 1+2+3 RL (Ours) 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.4})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ Advantage Shaping Linear Decay Weights 48.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.1})}$ 46.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.2})}$ 31.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 44.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.7})}$ 64.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.6})}$ 30.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ Linear Incline Weights 48.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.3})}$ 46.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.4})}$ 31.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 44.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.0})}$ 65.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.8})}$ 30.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.4})}$ Uniform Weights (Baseline GRPO) 47.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.6})}$ 45.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 31.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 43.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.1})}$ 64.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.6})}$ 29.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ Inverse Parabolic Weights 46.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.1})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.0})}$ 42.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.2})}$ 63.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.3})}$ 29.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.1})}$ TAS (Ours) 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.4})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$
5.1 Results
@@
As shown in Table 2, AVATAR substantially improves audio-visual reasoning over baselines like Qwen2.5-Omni (+4.9+4.9 on OmniBench, +3.0+3.0 on DailyOmni) and Ola-7B (+3.4+3.4 on DailyOmni, +3.2+3.2 on AV-Odyssey). Our stratified replay buffer addresses GRPO’s vanishing advantage issue, preventing stalling and overfitting and outperforming methods like AV-Reasoner and Omni-R1. TAS further enhances performance on benchmarks that require precise temporal alignment, such as AV-Odyssey. By upweighting tokens at the sequence start and end, TAS is designed to reinforce the critical reasoning steps such as early audio grounding and final cross-modal integration needed for such tasks.
@@
AVATAR demonstrates strong performance on general video understanding benchmarks like MVBench (+5.3+5.3 w/ Ola) and VideoMME (+4.5+4.5 w/ Qwen), and excels in complex reasoning (Table 3). It achieves significant gains on causal inference tasks such as Video-Holmes (+4.5+4.5) and MMVU (+5.4+5.4). These improvements stem from the stratified replay buffer ensuring exposure to hard examples and TAS guiding reasoning by emphasizing planning and synthesis phases. Unlike unstable statistical weighting methods (e.g., TW-GRPO [twgrpo] on TOMATO), TAS leverages temporal structure for a stable learning signal, yielding consistent top performance.
(a) GRPO
@@
Figure 4: Comparison of training dynamics between GRPO and AVATAR. GRPO (a) shows oscillatory and unstable accuracy reward progression, whereas AVATAR (b) demonstrates a smoother, more consistent learning trajectory.
Table 6: Ablation comparing the impact of SFT and AVATAR integration. Models trained with SFT achieve stronger cross-modal grounding and reward alignment, while removing SFT leads to significant degradation in both GRPO and AVATAR.
Setting Audio-Visual Benchmarks OmniBench DailyOmni AV-Odyssey Qwen2.5 Omni 44.2 44.0 29.8 GRPO (w/o SFT) 41.2 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-3.0})}$ 40.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-3.2})}$ 27.1 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-2.7})}$ GRPO (w/ SFT) 46.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 45.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.5})}$ 30.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.4})}$ AVATAR (w/o SFT) 43.8 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-0.4})}$ 43.5 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-0.5})}$ 28.9 $\scriptstyle{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(\textbf{-0.9})}$ AVATAR (w/ SFT) 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$
5.2 Ablation Studies
[…]
RQ2: What is the effect of each component in AVATAR?
Our component-wise ablation study (Table 4) isolates how each part of AVATAR addresses specific GRPO limitations. Introducing the replay buffer significantly boosts performance over baseline GRPO (e.g., +3.6+3.6 on OmniBench), confirming its effectiveness in tackling data inefficiency and the vanishing advantage problem by enabling off-policy learning, even with uniform credit assignment. Conversely, adding TAS to standard on-policy GRPO also yields gains (e.g., +4.1+4.1 on MMVU) over baseline GRPO, demonstrating its ability to improve performance solely by resolving uniform credit assignment. Combining both the replay buffer and TAS results in further improvements, as the buffer provides better data while TAS ensures effective learning from it. Finally, adding the hinting mechanism provides an additional performance lift, validating its role in helping the model escape challenging local optima.
RQ3: What is the effect of the training curriculum and advantage shaping?
Table 5 validates the effectiveness of both the staged training curriculum and the TAS strategy. The curriculum shows consistent stage-wise improvements: Stage 11 RL strengthens visual reasoning over SFT, Stage 22 adds cross-modal alignment, and Stage 33 refines fine-grained localization, yielding progressive gains in reasoning ability. The advantage shaping ablation compares several weighting schemes: uniform (GRPO baseline), linear decay (planning-focused), linear incline (synthesis-focused), inverse parabolic (middle-focused), and our parabolic TAS (planning and synthesis-focused). Results confirm that extreme emphasis on either phase or the middle yields inferior results, while TAS achieves the best overall performance by jointly amplifying early planning and late synthesis steps.
(a) Audio-Visual Reasoning
(b) Video Reasoning
Figure 5: Effect of reasoning sequence length on TAS performance. In both (a) audio-visual and (b) video reasoning benchmarks, TAS yields greater gains with longer reasoning sequences.
RQ4: Effect of SFT on Learning Stability.
Table 6 shows that incorporating SFT before GRPO provides essential structural grounding that improves convergence and generalization. Without SFT, models lack cross-modal alignment, resulting in unstable rewards and degraded performance across all benchmarks. Initializing GRPO from an SFT-trained policy yields more stable learning and moderate gains, while AVATAR achieves the highest improvements. Hence, SFT establishes a reliable foundation for stable advantage estimation and TAS-based reward shaping, enabling GRPO to refine rather than relearn multimodal reasoning.
RQ5: Impact of TAS across Reasoning Lengths.
As shown in Figure 5, TAS consistently improves performance across reasoning lengths, with gains amplifying as sequences grow longer. The advantage becomes more pronounced for extended sequences ($400+$ tokens), particularly on challenging benchmarks like MMVU and Video-Holmes. This highlights TAS’s ability to preserve credit assignment and stabilize optimization in long-horizon reasoning, where uniform credit tends to dilute reward signals.
6 Conclusion